# ALPCAHUS: Subspace Clustering for Heteroscedastic Data

### v2

@@ IV-A Ensemble Extension for ALPCAHUS
This section presents the ensemble algorithm with base clustering parameter $B$ to combine multiple trials for finding $\mathcal{C}$ in (15). Any $B>1$ leverages consensus clustering by forming an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ where
$\bm{W}_{i,j}=$ $\frac{1}{B}|\{\forall b\in\{1,\ldots,B\}\text{ such that }$
@@
$\bm{Z}^{\text{row}}=\operatorname*{arg\,min}_{\bm{Z}}\left\|\bm{W}-\bm{Z}\right\|_{\mathrm{F}}^{2}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\|\bm{W}_{i,:}\|_{0}=q\hskip 5.69054pt\forall i$ (23)
$\bm{Z}^{\text{col}}=\operatorname*{arg\,min}_{\bm{Z}}\left\|\bm{W}-\bm{Z}\right\|_{\mathrm{F}}^{2}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\|\bm{W}_{:,i}\|_{0}=q\hskip 5.69054pt\forall i.$ (24)
@@
Finally, spectral clustering is applied to $\frac{1}{2}(\bm{Z}^{\text{col}}+\bm{Z}^{\text{row}})$ to get the clusters from the results of the ensemble. One could select the base clusterings parameter $B=1$ to reduce to one trial of the optimization problem (15). Alg. ALPCAHUS summarizes the procedure for subpace clustering with optional ensemble learning. For Julia code implementations, refer to https://github.com/javiersc1/ALPCAHUS.
IV-B Rank Estimation
@@
In subspace clustering, some algorithms like SSC do not require subspace dimension to be known or estimated, whereas others like KSS require this. For the group of algorithms that require this parameter to be estimated, there is great interest in adaptive methods that can learn the subspace dimension. In recent work, [26] proposed using an eigengap heuristic on the sample covariance matrix of each cluster, i.e., $S_{k}=\frac{1}{|\mathcal{C}_{k}|}\bm{Y}_{k}\bm{Y}_{k}^{T}$, to estimate dimension. This means calculating the following
$\hat{d}_{k}=\operatorname*{arg\,max}_{i}|\lambda_{i}(S_{k})-\lambda_{i+1}(S_{k})|$ (25)
@@
where $\lambda_{i}(S_{k})$ denotes the $i$th eigenvalue of $S_{k}$ assuming $\lambda_{1}\geq\ldots\geq\lambda_{D}$. For heteroscedastic data, the eigengap heuristic can break down, making it challenging to determine rank, as shown in Sec. V, Fig. 5. In recent work, [34] developed a parallel analysis algorithm to estimate the rank of a matrix that is consistently shown to work well in the heteroscedastic regime if the data comes from one subspace only. It works by creating an i.i.d. Bernoulli $(p=0.5)$ matrix denoted as $\bm{M}$ and analyzing the singular values of $\bm{M}\odot\bm{Y}$ for a matrix of interest $\bm{Y}$. One can distinguish what singular values are associated with the signal and noise component of the data through this process. We generalize this line of work and apply it to the union of subspace setting by starting off over-parameterized and adaptively shrinking subspace bases as follows:
@@
This is done for each estimated cluster $\bm{Y}_{k}$ over $R$ random trials where $1\leq R\ll T_{2}=\text{ALPCAHUS maximum total iterations}$. Here $\sigma_{d+1}(\bm{Y}_{k})$ denotes the $d+1$th singular value of $\bm{Y}_{k}$. This process is repeated for all cluster subsets $\{\bm{Y}_{1},\ldots,\bm{Y}_{K}\}$ after the cluster reassignment update in (17). To reduce computation, this dimension estimation is performed sparingly every $10\%$ of ALPCAHUS iterations.
IV-C Cluster Initialization
For the non-ensemble version of ALPCAHUS ($B=1$), it previously remained to be seen whether there exists an initialization scheme that performs better in expectation than random cluster assignment in the heteroscedastic regime. In recent work, [26] proposes a thresholding inner-product based spectral initialization method (TIPS), designed for homoscedastic data to be used with KSS, where an affinity matrix is generated by
$\bm{W}_{ij}=1\text{ if }|\langle\bm{y}_{i},\bm{y}_{j}\rangle|\geq\tau\text{ and }i\neq j$ (28)
given a thresholding parameter $\tau>0$ and zero otherwise. Then, the cluster assignments $\mathcal{C}=\{c_{1},\ldots,c_{N}\}$ are calculated by applying the spectral clustering method on $\bm{W}$. Recall that $\bm{y}_{i}=\bm{x}_{i}+\bm{\epsilon}_{i}$. Upon closer analysis, this metric in expectation, ignoring absolute value, gives
@@ IV-C Cluster Initialization
IV-D ALPCAHUS Convergence
ALPCAHUS with a single cluster ($K=1$) has a sequence of cost function values that provably converges to local minima, since the cost function and algorithm simplifies to LR-ALPCAH (14) that has convergence guarantees shown in [17, Thm. 2]. In the multi-cluster setting ($K>1$), Thm. 1 shows that one can guarantee convergence of the sequence of cost function values in (15) by ensuring two things: that the noise variances are kept above some small threshold value, and that the cluster assignment criteria makes assignment changes only if the cost function strictly decreases. The argument in Thm. 1 below is a natural extension of that in the original k-subspaces paper [25, Thm. 7]. The Appendix provides an experimental result, Fig. 9, that corroborates this theorem.
Theorem 1.
Consider the non-ensemble version of the ALPCAHUS algorithm. Assume a stopping criteria where the loop over $t_{2}$ in the algorithm continues until there is an iterate where the cost function does not decrease. Assume a noise variance threshold parameter $\alpha\in\mathbb{R}>0$ that lower bounds all $\nu_{i}$ and a cluster assignment criteria that accepts changes only if there is a cluster reassignments of points that strictly decrease the cost function, as expressed in (17). Then, the ALPCAHUS algorithm generates a sequence of cost function values that converges to a fixed point.
Proof.
To prove that the sequence of cost function values converges, one needs to assess whether each step decreases the cost function, and whether the algorithm terminates at a fixed point where the variables are no longer updated. Each step consists of two sub-steps: first, the subspace estimation sub-step where LR-ALPCAH is used, and second, the clustering sub-step where projection onto the estimated subspaces is done to cluster with respect to the smallest residual. For the subspace estimation sub-step, with cluster assignments fixed, it is known that LR-ALPCAH has a sequence of cost function values that will converge according to guarantees given in [17, Thm. 2]. For the clustering sub-step, given our stopping criteria, once the clusters do not change, the algorithm has a sequence of cost function values that reaches a fixed point. That leaves only the case where the clusters do change, which will be the focus in the remainder of the proof. Let $\bm{Y}_{k}^{(t)},\bm{L}_{k}^{(t)},\bm{R}_{k}^{(t)},\bm{\Pi}_{k}^{(t)}$ be the optimization variables that produce the cost
$\text{cost}^{(t_{1},t_{2})}=$ $\sum_{k}\frac{1}{2}\|(\bm{Y}_{k}^{(t_{2})}-\bm{L}_{k}^{(t_{1})}[\bm{R}_{k}^{(t_{1})}]^{T})[\bm{\Pi}_{k}^{(t_{1})}]^{-\frac{1}{2}}\|_{F}^{2}$
$+\frac{D}{2}\log{|\bm{\Pi}_{k}^{(t_{1})}|}$ (31)
where $t_{1}$ is associated with the variables in the subspace estimation sub-step and $t_{2}$ is associated with the variables in the clustering sub-step. Recall that the clustering is produced by solving
$c_{i}=\operatorname*{arg\,min}_{k}\|\bm{y}_{i}-\bm{U}_{k}^{(t_{1})}[\bm{U}_{k}^{(t_{1})}]^{T}\bm{y}_{i}\|_{2}^{2}$ (32)
for all data points $y_{i}$, where $\bm{U}_{k}^{(t_{1})}$ is found by performing an SVD on $\bm{L}_{k}^{(t_{1})}$ or by $\bm{U}_{k}^{(t_{1})}=\bm{L}_{k}^{(t_{1})}[\bm{L}_{k}^{(t_{1})}]^{\dagger}$. Under this setting, there is at least one $\bm{y}_{i}$ with label $c_{i}^{(t_{2}+1)}\neq c_{i}^{(t_{2})}$. For those points, denote the previous cluster assignment as $\hat{k}$. Due to the cluster assignment criteria, it is know that
$\min_{k}\|\bm{y}_{i}-\bm{U}_{k}^{(t_{1}+1)}[\bm{U}_{k}^{(t_{1}+1)}]^{T}\bm{y}_{i}\|_{2}^{2}$
$\leq\|\bm{y}_{i}-\bm{U}_{\hat{k}}^{(t_{1}+1)}[\bm{U}_{\hat{k}}^{(t_{1}+1)}]^{T}\bm{y}_{i}\|_{2}^{2}$ (33)
$\implies(\bm{y}_{i}\in\mathcal{C}_{\hat{k}}^{(t_{2})}\implies\bm{y}_{i}\in\mathcal{C}_{k}^{(t_{2}+1)}).$ (34)
It is easy to see that $\bm{X}_{k}=\bm{L}_{k}\bm{R}_{k}^{T}=\bm{U}_{k}\bm{U}_{k}^{T}\bm{Y}_{k}$ since the projection of $\bm{Y}_{k}$ onto $\mathcal{R}(\bm{U}_{k})$ is $\bm{X}_{k}$. The cost can be expressed as
$\text{cost}^{(t_{1},t_{2})}$ $=\sum_{k}\frac{1}{2}\|(\bm{Y}_{k}^{(t_{2})}-\bm{X}_{k}^{(t_{1})})[\bm{\Pi}_{k}^{(t_{1})}]^{-1/2}\|$
$+\frac{D}{2}\log{|\bm{\Pi}_{k}^{(t_{1})}|}.$ (35)
Then, the following observation is made regarding the cost function written in (35).
$\text{cost}^{(t_{1},t_{2})}$ $=\sum_{k}\sum_{i\in\mathcal{C}_{k}^{(t_{2})}}\frac{1}{\sqrt{\nu_{i}^{(t_{1})}}}\|[\bm{Y}_{k}]_{i}-[\bm{X}_{k}]_{i}^{(t_{1})}\|_{2}^{2}$
$+\frac{D}{2}\log{|\bm{\Pi}_{k}^{(t_{1})}|}$ (36)
$\geq\sum_{k}\sum_{i\in\mathcal{C}_{k}^{(t_{2})}}\frac{1}{\sqrt{\nu_{i}^{(t_{1}+1)}}}\|[\bm{Y}_{k}]_{i}-[\bm{X}_{k}]_{i}^{(t_{1}+1)}\|_{2}^{2}$
$+\frac{D}{2}\log{|\bm{\Pi}_{k}^{(t_{1}+1)}|}=\text{cost}^{(t_{1}+1,t_{2})}$ (37)
$\geq\sum_{k}\sum_{i\in\mathcal{C}_{k}^{(t_{2}+1)}}\frac{1}{\sqrt{\nu_{i}^{(t_{1}+1)}}}\|[\bm{Y}_{k}]_{i}-[\bm{X}_{k}]_{i}^{(t_{1}+1)}\|_{2}^{2}$
$+\frac{D}{2}\log{|\bm{\Pi}_{k}^{(t_{1}+1)}|}=\text{cost}^{(t_{1}+1,t_{2}+1)}$ (38)
Here, (IV-D) is due to the subspace step not increasing the cost as shown in [17, Thm. 2], and (IV-D) is due to the argument above in (IV-D). What has been shown is that
$\text{cost}^{(t_{1},t_{2})}\geq\text{cost}^{(t_{1}+1,t_{2})}\geq\text{cost}^{(t_{1}+1,t_{2}+1)}.$ (39)
Therefore, as the cluster assignments and subspace bases are updated, then the cost will either decrease or remain the same. Once the cost remains the same, then the algorithm has a sequence of cost function values that converges to a fixed point with locally optimal cluster labels. ∎
@@ V-A1 Experimental Setup
A synthetic dataset is generated consisting of $K=2$ clusters each of dimension $d=3$ derived from random subspaces in $D=100$ dimensional ambient space. Each cluster consisted of two data groups with group 1 containing $N_{1}=6$ samples, to explore the data constrained regime, with noise variance $\nu_{1}=0.1$ per cluster. For group 2, the $N_{2}$ samples and noise variance $\nu_{2}$ are varied. Cross validation is done for hyperparameters in any algorithms that require it with a separate training set. Both the estimated clusters from each clustering algorithm and the known clusters are necessary to calculate clustering error. Additionally, the problem of label permutations is dealt by using the Hungarian algorithm [35]. More concretely, clustering error is defined as
clustering $\text{clustering error}=\frac{100}{N}(1-\max_{\pi}\sum_{i,j}\bm{Q}_{\pi(ij)}^{\text{out}}\bm{Q}_{ij}^{\text{true}})$ (40)
where $\pi$ is a permutation of cluster labels, and $\bm{Q}^{\text{out}}$ and $\bm{Q}^{\text{true}}$ are output labels and ground truth labelings of the data where the $(i,j)$th entry is one if point $j$ belongs to cluster $i$ and zero otherwise.
The average clustering error is computed out of 100 trials where each trial had different noise, basis coefficients, and subspace basis realizations for the clusters. Various subspace clustering algorithms are compared such as $K$-Subspaces (KSS) [14], Ensemble $K$-Subspaces (EKSS) [16], Subspace Clustering via Thresholding (TSC) [36], and Doubly Stochastic Sparse Subspace Clustering (ADSSC) [37]. Applying general non-subspace clustering methods such as $K$-means [38] resulted in poor clustering error so results of these methods are generally not shown.

### v3

@@ IV-A Ensemble Extension for ALPCAHUS
This section presents the ensemble algorithm with base clustering parameter $B$ to combine multiple trials for finding $\mathcal{C}$ in (19). ALPCAHUS with $B>1$ leverages consensus clustering by forming an affinity matrix $\bm{A}\in\mathbb{R}^{N\times N}$ where
$\bm{A}_{i,j}=$ $\frac{1}{B}|\{\forall b\in\{1,\ldots,B\}\text{ such that }$
@@
$\bm{Z}^{\text{row}}=\operatorname*{arg\,min}_{\bm{Z}}\left\|\bm{A}-\bm{Z}\right\|_{\mathrm{F}}^{2}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\|\bm{Z}_{i,:}\|_{0}=q\hskip 5.69054pt\forall i$ (31)
$\bm{Z}^{\text{col}}=\operatorname*{arg\,min}_{\bm{Z}}\left\|\bm{A}-\bm{Z}\right\|_{\mathrm{F}}^{2}\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\|\bm{Z}_{:,i}\|_{0}=q\hskip 5.69054pt\forall i.$ (32)
Finally, spectral clustering is applied to $\bm{W}=\frac{1}{2}(\bm{Z}^{\text{col}}+\bm{Z}^{\text{row}})$ to get the clusters from the results of the ensemble. This spectral clustering operation assumes a balanced cluster distribution that may not hold in practice. All of our included experiments contain balanced cluster sizes; however, Sec. VIII-B provides discussion of this limitation and possible remediation strategies.
The $q$ threshold parameter can be set by cross validation, which is what is done in the experiments shown in Sec. V, or other techniques can be used. For example, Ref. [l2graph_subspace_clustering] creates a sparse L2 graph through hard thresholding to retain the top $q=d_{k}$ values where $d_{k}$ is the known subspace dimension. Thus, $q$ can be tied to the true or estimated subspace dimension instead. Sec. VIII-C provides more discussion on ALPCAHUS hyperparameters including recommended values.
@@
In this more general ensemble framework, one could select $B=1$ to reduce to one trial of the optimization problem in (19) which is the non-ensemble version of ALPCAHUS. Alg. ALPCAHUS summarizes the procedure for subpace clustering with optional ensemble learning. For Julia code implementations, refer to https://github.com/javiersc1/ALPCAHUS.
IV-B ALPCAHUS Convergence
In the single cluster setting, ALPCAHUS with $K=1$ has a sequence of cost function values $f_{1}(\bm{L}_{1},\bm{R}_{1},\bm{\Pi}_{1};\bm{Y}_{1})$ that provably converges to local minima, since the cost function and algorithm simplify to LR-ALPCAH (17) that has convergence guarantees as proven in Ref. [alpcah_journal, Thm. 2]. In the multi-cluster setting with $K>1$, Thm. 1 below shows that the sequence of cost function values produced by $f(\mathcal{L},\mathcal{R},\bm{\Pi},\mathcal{C})$ in (19) converges asympototically. The argument in Thm. 1 below is a natural extension of that in the original k-subspaces paper [kplane, Thm. 7] with greater mathematical exposé. Sec. VIII-A provides an experimental result that corroborates this theorem.
Theorem 1.
Consider the ALPCAHUS cost function $f(\mathcal{L},\mathcal{R},\bm{\Pi},\mathcal{C})$ in (19). Assume a noise variance threshold parameter $\alpha\in\mathbb{R}>0$ that lower bounds all $\nu_{i}$, and the cluster assignment criteria in (29) that accepts changes only if there is a cluster reassignment of points that strictly decreases the cost function $f(\cdot)$, as expressed in (29). Then, (19) generates a sequence of cost function values that converges asymptotically. Furthermore, the algorithm terminates in finite iterations due to stopping criteria in line (10) of Alg. ALPCAHUS that checks whether clusters have changed from the previous iteration.
Proof.
To prove that the sequence of cost function values converges, we show that each step decreases the cost function. We also show that the algorithm terminates at a point where the variables stop changing. Each step consists of two sub-steps: first, the subspace estimation sub-step where LR-ALPCAH is used, and second, the clustering sub-step where projection onto the estimated subspaces is done. Let $t_{1}$ be the iteration variable associated with the variables in the subspace estimation sub-step and $t_{2}$ be the iteration variable associated with the variables in the clustering sub-step. Define $\bm{X}_{k}=\bm{L}_{k}\bm{R}_{k}^{\prime}$ and bound the cost by
$f(\mathcal{L}^{(t_{1})},\mathcal{R}^{(t_{1})},\bm{\Pi}^{(t_{1})},\mathcal{C}^{(t_{2})})=$
$\frac{1}{2}\sum_{k}\sum_{i\in\mathcal{C}_{k}^{(t_{2})}}\frac{1}{\nu_{i}^{(t_{1})}}\left\|[\bm{Y}_{k}]_{i}-[\bm{X}_{k}]_{i}^{(t_{1})}\right\|_{2}^{2}+D\log{\nu_{i}^{(t_{1})}}$ (33)
$\geq\frac{1}{2}\sum_{k}\sum_{i\in\mathcal{C}_{k}^{(t_{2})}}\frac{1}{\nu_{i}^{(t_{1}+1)}}\left\|[\bm{Y}_{k}]_{i}-[\bm{X}_{k}]_{i}^{(t_{1}+1)}\right\|_{2}^{2}$
$+D\log\nu_{i}^{(t_{1}+1)}=f(\mathcal{L}^{(t_{1}+1)},\mathcal{R}^{(t_{1}+1)},\bm{\Pi}^{(t_{1}+1)},\mathcal{C}^{(t_{2})})$ (34)
$\geq\frac{1}{2}\sum_{k}\sum_{i\in\mathcal{C}_{k}^{(t_{2}+1)}}\frac{1}{\nu_{i}^{(t_{1}+1)}}\left\|[\bm{Y}_{k}]_{i}-[\bm{X}_{k}]_{i}^{(t_{1}+1)}\right\|_{2}^{2}$
$+D\log\nu_{i}^{(t_{1}+1)}=f(\mathcal{L}^{(t_{1}+1)},\mathcal{R}^{(t_{1}+1)},\bm{\Pi}^{(t_{1}+1)},\mathcal{C}^{(t_{2}+1)}).$ (35)
Here, (33) $\geq$ (IV-B) is due to the subspace step not increasing the cost for each $f_{k}$ term as shown in Ref. [alpcah_journal, Thm. 2], and (IV-B) $\geq$ (35) is due to the cluster reassignment criteria in (29) only updating the cluster label that decreases (25). What has been shown is that
$f^{(t_{1},t_{2})}\geq f^{(t_{1}+1,t_{2})}\geq f^{(t_{1}+1,t_{2}+1)}$ (36)
where $f^{(t_{1},t_{2})}=f(\mathcal{L}^{(t_{1})},\mathcal{R}^{(t_{1})},\bm{\Pi}^{(t_{1})},\mathcal{C}^{(t_{2})})$. Thus, as the cluster assignments and subspace bases are updated, the cost is non-increasing. Since the cost function values are non-increasing at every step, bounded below, then the sequence of cost function values produced by (19) converges asymptotically. Furthermore, the stopping criteria in line (10) ensures that the algorithm terminates when all clusters do not change. Because there are only a finite number of ways to assign cluster labels, Alg. ALPCAHUS will stop after a finite number of iterations. ∎
IV-C Rank Estimation
@@
In subspace clustering, some algorithms like SSC do not require subspace dimension to be known or estimated, whereas others like KSS require it. For the group of algorithms that require this parameter, there is great interest in adaptive methods that can learn the subspace dimension. In recent work, Ref. [kss_convergence] proposed using an eigengap heuristic on the sample covariance matrix of each cluster to estimate dimension. Mathematically, this means calculating the following
$\bm{S}_{k}$ $=\frac{1}{|\mathcal{C}_{k}|}\bm{Y}_{k}\bm{Y}_{k}^{\prime}$ (37)
$\hat{d}_{k}$ $=\operatorname*{arg\,max}_{i}|\lambda_{i}(\bm{S}_{k})-\lambda_{i+1}(\bm{S}_{k})|$ (38)
@@
where $\lambda_{i}(S_{k})$ denotes the $i$th eigenvalue of $S_{k}$ assuming $\lambda_{1}\geq\ldots\geq\lambda_{D}$. For heteroscedastic data, the eigengap heuristic can break down, making it challenging to determine rank, as shown in Sec. V, Fig. 5. In recent work, Ref. [signFlips] developed a parallel analysis algorithm to estimate the rank of a matrix that is consistently shown to work well in the heteroscedastic regime if the data comes from a single subspace only. It works by creating an i.i.d. Bernoulli $(p=0.5)$ matrix denoted as $\bm{M}$ and analyzing the singular values of $\bm{M}\odot\bm{Y}$ for a matrix of data samples $\bm{Y}$. One can distinguish what singular values are associated with the signal and noise component of the data through this process. We generalize this line of work and apply it to the union of subspace setting by starting off over-parameterized and adaptively shrinking the bases by
@@
This is done for each estimated cluster $\bm{Y}_{k}$ over $R$ random trials where $1\leq R\ll T_{2}$, and $T_{2}$ denotes the maximum ALPCAHUS iterations. Here, $\sigma_{d+1}(\bm{Y}_{k})$ denotes the $d+1$ singular value of $\bm{Y}_{k}$. This process is repeated for all cluster subsets $\{\bm{Y}_{1},\ldots,\bm{Y}_{K}\}$ after the cluster reassignment update in (29). To reduce computation, this dimension estimation is performed sparingly in the ALPCAHUS method.
IV-D Cluster Initialization
For the non-ensemble version of ALPCAHUS ($B=1$), it previously remained to be seen whether there exists an initialization scheme that performs better than random cluster assignment in the heteroscedastic regime. In recent work, Ref. [kss_convergence] proposed a thresholding inner-product based spectral initialization method (TIPS), designed for homoscedastic data to be used with KSS, where an affinity matrix is generated by $\bm{W}_{ij}=1\text{ if }|\langle\bm{y}_{i},\bm{y}_{j}\rangle|\geq\tau$ and $i\neq j$ given a thresholding parameter $\tau>0$. Instead, in this work, we generate a fully connected $\bm{W}$ by creating
$\bm{W}_{ij}=|\langle\bm{y}_{i},\bm{y}_{j}\rangle|\text{ if }i\neq j,\hskip 2.84526pt0\text{ otherwise},$ (41)
and hard threshold to retain the top $q$ edges by solving (11) so that $\tau=q$ for simplicity. Then, the cluster assignments $\mathcal{C}=\{c_{1},\ldots,c_{N}\}$ are calculated by applying the spectral clustering method on the thresholded $\bm{W}$. Recall that $\bm{y}_{i}=\bm{x}_{i}+\bm{\epsilon}_{i}$. Upon closer analysis, this metric in expectation, ignoring absolute value, gives
@@ V-A1 Experimental Setup
A synthetic dataset is generated consisting of $K=2$ clusters each of dimension $d=3$ derived from random subspaces in $D=100$ dimensional ambient space. Each cluster consisted of two data groups with group 1 containing $N_{1}=6$ samples, to explore the data constrained regime, with noise variance $\nu_{1}=0.1$ per cluster. For group 2, the $N_{2}$ samples and noise variance $\nu_{2}$ are varied. Cross validation is done for hyperparameters in any algorithms that require it with a separate training set. For a more detailed description of ALPCAHUS parameters, including recommended default values, see Sec. VIII-C. Both the estimated clusters from each clustering algorithm and the known clusters are necessary to calculate clustering error. Further, the problem of label permutations is dealt with by using the Hungarian algorithm [hungarian]. More concretely, clustering error (%) is defined as
$\text{clustering error}=\frac{100}{N}\left(1-\max_{\pi}\sum_{i,j}\bm{Q}_{\pi(ij)}^{\text{out}}\bm{Q}_{ij}^{\text{true}}\right)$ (44)
where $\pi$ is a permutation of cluster labels found by applying the Hungarian algorithm, and $\bm{Q}^{\text{out}}$ and $\bm{Q}^{\text{true}}$ are output labels and ground truth labelings of the data where the $(i,j)$th entry is one if point $j$ belongs to cluster $i$ and zero otherwise.
The average clustering error is computed out of 100 trials where each trial had different noise, basis coefficients, and subspace basis realizations for the clusters. Various subspace clustering algorithms are included such as $K$-Subspaces (KSS) [k-subspaces], Ensemble $K$-Subspaces (EKSS) [ekss], Subspace Clustering via Thresholding (TSC) [tsc], and Doubly Stochastic Sparse Subspace Clustering (ADSSC) [adssc] which is a modern variant of SSC. Applying general non-subspace clustering methods such as $K$-means [kmeans] resulted in poor clustering error so results of these methods are not shown in general.